OpenStudy (adunb8):

this is linear algebra & differential equation about Homogenous Equation question... convert to separable eqs: v=y/x 1.) xyy' = 2y^2-x^2

OpenStudy (lgbasallote):

change y' to dy/dx \[\implies (xy)\frac{dy}{dx} = 2y^2 - x^2\] cross multiply.. \[\implies (xy)dy = (2y^2 - x^2)dx\] put everything on one side... \[\implies (2y^2- x^2)dx - (xy)dy = 0\] let y = vx

OpenStudy (lgbasallote):

so dy = vdx + xdv

OpenStudy (lgbasallote):

\[\implies (2v^2 x^2 - x^2)dx - (vx^2)(vdx + xdv) = 0\] multiply the thingies... \[\implies 2v^2 x^2 dx - x^2 dx - v^2 x^2 dx - vx^3 dv = 0\] are you still following?

OpenStudy (lgbasallote):

i think @adunb8 went offline o.O

OpenStudy (adunb8):

yea im on sorry

OpenStudy (lgbasallote):

so you get what im doing so far?

OpenStudy (adunb8):

yes! omg! but teacher wants me to determine if its separable or not then use the v=y/x so how do determine it is separable eqn?

OpenStudy (lgbasallote):

all homogeneous are separable in the later stage...which im going to do now

OpenStudy (adunb8):

okay but how do i determine that on the test because on the test he will only say use whatever your learned so far and he wont specify which to use

OpenStudy (lgbasallote):

first combine like terms... \[\implies v^2 x^2 dx -x^2 dx - vx^3 dv\] factor out \[\implies x^2dx(v^2 - 1) - vx^3dv\] do you see the separable thingy now?

OpenStudy (lgbasallote):

i can tell you how to know if it's homogeneous...

OpenStudy (adunb8):

okay=)

OpenStudy (lgbasallote):

you look at the degree...when the degree are all the same...it's homogeneous for example (xy) <--degree 2 (2y^2) <--degree 2 (x^2) <--degree 2 therefore, it's homogeneous

OpenStudy (adunb8):

why is (xy) degree of 2?

OpenStudy (lgbasallote):

x^1 y^1 <--1 + 1 = 2

OpenStudy (adunb8):

oh i see sorry i dont really see things right away...

OpenStudy (lgbasallote):

note: ALL terms have to be SAME degree for it to be homo

OpenStudy (adunb8):

i see!!

OpenStudy (lgbasallote):

for exact de \[\frac{\partial M}{\partial y} = \frac{\partial N}{\partial x}\]

OpenStudy (lgbasallote):

for linear DE \[\frac{dy}{dx} + P(x) y = Q(x)\]

OpenStudy (adunb8):

what about for separable?

OpenStudy (lgbasallote):

i am not aware of any mark for it. i think it's just trial and error

OpenStudy (adunb8):

oh okay... i guess that really has to be trial and error

OpenStudy (lgbasallote):

but usually it's in the form \[f(x)dx + g(y)dy = 0\]

OpenStudy (lgbasallote):

that means dx only has function of x and dy only has function of y

OpenStudy (lgbasallote):

however, there are times, when you can manipulate an equation into becoming separable

OpenStudy (adunb8):

okay so after i plug into the v= y/x formula then what do i do in order to solve the final solution?

OpenStudy (lgbasallote):

you dont do v = y/x yet like i said your FIRST substitution is y = vx

OpenStudy (lgbasallote):

you do v = y/x in the back subsitution... AFTER you get the solution

OpenStudy (adunb8):

okay so after y=vx i plug back into the original solution?

OpenStudy (lgbasallote):

are we talking about the first step?

OpenStudy (adunb8):

im talking about here you said above couple of replies earlier. first combine like terms... ⟹v2x2dx−x2dx−vx3dv factor out ⟹x2dx(v2−1)−vx3dv do you see the separable thingy now?

OpenStudy (lgbasallote):

oh.. now you do variable separable

OpenStudy (lgbasallote):

you solve for the solution

OpenStudy (adunb8):

i think i need to solve for y but i see that you changed all of them to y =vx

OpenStudy (lgbasallote):

you dont solve for anything yet....

OpenStudy (lgbasallote):

you just solve for the GENERAL SOLUTION

OpenStudy (adunb8):

how do i do that from here? use by doing integral steps?

OpenStudy (lgbasallote):

like i said...variable separable

OpenStudy (adunb8):

i dont really understand what you mean by variable separable..... sorry =(

OpenStudy (lgbasallote):

\[x^2 dx(v^2 - 1) - vx^3 dv\] \[\implies \frac{x^2dx}{x^3} - \frac{vdv}{v^2 - 1}\] that's what variable separation means

OpenStudy (adunb8):

oh isee okay !

OpenStudy (lgbasallote):

so solve the general solution now

OpenStudy (adunb8):

by doing the integral right? and using u subsitution?

OpenStudy (lgbasallote):

yes

OpenStudy (adunb8):

thank you so so much you are my life saver every time! !! =)

OpenStudy (lgbasallote):

welcome

OpenStudy (lgbasallote):

after you integrate..change v into y/x

OpenStudy (lgbasallote):

then that would be the final answer

OpenStudy (adunb8):

right back to original!

OpenStudy (lgbasallote):

yep